# Entropy-driven entanglement forging

### v1

@@ IV.3 Number of CNOT gates in EDEF NSM simulations
Figure 8 compares the number of CNOT gates required in the NSM simulation of 28Ne and 60Ti in the case of regular ADAPT-VQE and EDEF with one or two layers. Since the latter cases involve several circuits, Fig. 8 selects the one with maximum number of CNOT gates. The top panel shows that, for 28Ne, the number of CNOTs in the 24-qubit circuit (ADAPT-VQE, solid red line) is drastically reduced in one order of magnitude throughout the whole optimization by EDEF with one layer using 12 qubits (1-cut EDEF, dashed blue line). The 6-qubit EDEF with two layers (2-cut EDEF, dotted-dashed yellow line) reduces the number of CNOTs in an additional order of magnitude. Consistently, the bottom panel of Fig. 8 indicates that for 60Ti the number of CNOTs is smaller in the 2-cut EDEF, which uses 10-qubit circuits, than in the 1-cut EDEF involving 20 qubits. These results highlight that for these systems, our EDEF approach, in addition to having better performance than the regular ADAPT-VQE as shown in Sec. IV.2, is also better suited to be implemented in near-future quantum devices.
@@
The reason for the reduction of CNOT gates with EDEF is twofold. First, each additional EDEF cut makes circuits much shallower because at each iteration an operator is added to a single circuit, distributing the CNOT gates among all circuits rather than accumulating them in a single one. Moreover, circuits with fewer qubits constrain the fermionic operators in the ADAPT-VQE pool to be more local, requiring less CNOT gates in the Jordan-Wigner mapping.
@@
Figure 8: Number of CNOT gates for the circuit with maximum number for the simulation of 28Ne (top) and 60Ti (bottom) as a function of the ADAPT-VQE iterations. Results for regular ADAPT-VQE (0 cuts, solid red line), EDEF with one layer (1 cut, blue dashed line) and two layers (2 cuts, dashed-dotted yellow lines).
@@
Among EDEF circuits, both for 60Ti and 28Ne, the ones with most CNOT gates are those simulating the first, non-degnerate, product state in the Schmidt decomposition, $\lambda_{0}|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle$ in Eq. (14). This is because the ADAPT-VQE algorithm finds, in most iterations, largest gradients for operators added to the first, $J_{z}=0$ proton-neutron product state. This is expected from the Schmidt decomposition, as the first singular value is much larger than the rest and should contribute most to lowering the ground-state energy. Moreover, we find that ADAPT-VQE chooses the same circuit for many contiguous iterations, showing plateaus that indicate a constant number of CNOTs in all but one circuit during various iterations. More details on the specifics of the ADAPT-VQE optimization with EDEF are given in Appendix C, which also gives additional details about the number of CNOT gates in each of the circuits that appear in the EDEF approach.
@@ V Summary and outlook
We present a novel procedure, entropy-driven entanglement forging (EDEF), that optimizes the solution of many-body problems in quantum computers using variational quantum algorithms (VQAs). Our approach is based on entanglement forging, which reduces the number of qubits and quantum gates needed to solve the problem. Crucially, in order to decide on the corresponding partitions and coefficients, EDEF adds criteria associated with the entropy between partitions of the system.
@@
With this technique, we have successfully simulated ground states of Fermi-Hubbard (FH) lattices with various central hoppings and interactions, and of the isotopes 28Ne and 60Ti calculated with the nuclear shell model (NSM), in both cases using EDEF with ADAPT-VQE. For these systems, EDEF exploits the low entanglement between the left and right parts of FH lattices or between protons and neutrons in the NSM, and this knowledge is combined with the singular-value degeneracy in the Schmidt decomposition of the ground states. Our EDEF simulations with one layer of EDEF – separating the system in two bipartitions – only need half as many qubits as the regular ADAPT-VQE, and they describe nuclei with an order of magnitude fewer CNOT gates. Nuclei simulated with two EDEF layers – each bipartition additionally separated into two parts corresponding to higher- and lower-energy single-particle states – need a quarter of the regular ADAPT-VQE qubits and an additional order of magnitude fewer CNOTs.
@@
In terms of performance, in FH lattices EDEF converges better than regular ADAPT-VQE for weak central hopping, and it still works well as long as long as the central hopping equals the other hoppings, $t_{m}\approx t$. In the NSM, for 28Ne, one layer of EDEF using 8 circuits is more efficient than the regular ADAPT-VQE, while in 26Ne EDEF performs slightly worse due to the larger proton-neutron entanglement. Furthermore, EDEF allows us to simulate 60Ti, which is beyond our capabilities with the standard ADAPT-VQE. For this nucleus, the two layer EDEF reaches a lower infidelity than the 1-cut EDEF.
@@
In summary, EDEF is ideal to adapting to qubit and CNOT number limitations present in current intermediate-scale noisy devices. In particular, EDEF allows one to adjust the algorithm to produce shallower circuits mitigating the impact of errors. This reduction can be further improved by the adaptability of the trained parameters expected for a variational algorithm. Furthermore, by reducing the number of qubits EDEF allows one us to use smaller devices with lower error rates. In principle, these advantages can be exploited further by identifying low-entanglement partitions in successive EDEF layers. Beyond the first applications studied in this work, EDEF can optimize other algorithms, as well as other many-body systems for which low-entropy subsystems are identified.
@@ Acknowledgements.
This work is financially supported by MCIN/AEI/10.13039/5011 00011033 from the following grants: PID2020-118758GB-I00, PID2021-127890NB-I00, RYC-2017-22781 and RYC2018-026072 through the “Ramón y Cajal” program funded by FSE “El FSE invierte en tu futuro”, CNS2022-135529 and CNS2022-135716 funded by the “European Union NextGenerationEU/PRTR”, and CEX2019-000918-M to the “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences; and by the Generalitat de Catalunya, grants 2021SGR01095 and 2021SGR00907.
@@ A.1 First layer of entanglement forging
For the nuclear shell model and for all isotopes considered, with one layer of entanglement forging we separate the proton and neutrons degrees of freedom as shown in Fig. 4 and thus reduce by half the number of qubits of the simulating circuits. As observed in Fig. 5, the nuclear state can be, to a very good approximation, be approximated using the following truncation in the Schmidt decomposition
@@
With one circuit, for example, we simulate the state $|p_{-2}\rangle$, which corresponds to a superposition of Slater determinants containing only proton orbitals and with expected third component of total angular momentum $\langle p_{-2}|J_{z}|p_{-2}\rangle=-2$. We need to ensure these states are orthogonal. All states that have different $J_{z}$ expected value in each partition are already orthogonal, and will continue to be considering the operators in our pool, but this is not the case for $|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle$ and $|p_{0}\rangle\otimes|n_{0}\rangle$. For these, we start with different (orthogonal) Slater determinants, $|\psi_{0}\rangle$, $|\tilde{\psi}_{0}\rangle$, and then apply the same unitaries to both states to keep them orthogonal, that is, $\langle\tilde{\psi}_{0}|e^{-i\theta T}e^{i\theta T}|\psi_{0}\rangle=\langle\tilde{\psi}_{0}|\psi_{0}\rangle=0$.
At each iteration the gradients are computed for each of the terms in Eqs. (13) and (14) separately, except for $|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle$ and $|p_{0}\rangle\otimes|n_{0}\rangle$, as they both have the same parameters. In this case $|\psi_{c}\rangle$ in Eq. (10) represents $\lambda_{0}|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle+\lambda_{1}|p_{0}\rangle\otimes|n_{0}\rangle$. Once all the gradients are computed, the largest one determines which operator from the pool is chosen and in which circuit is implemented. To compute the energy, we extract the statevector from each circuit, make the corresponding tensor products between proton and neutron states, add them up as in Eq. (14), and compute the expected value of the Hamiltonian with the obtained statevector. The energy computation in an actual quantum computer could be implemented as proposed in [25], with the exception of crossed statistics in the two non-orthogonal states.
@@ A.2 Second layer of entanglement forging
To add only few more parameters to the ansatz we decompose each of the previous statevectors with well defined spin and isospin into only two terms with different particle distributions. For the first term in Eq. 14, with $S_{z}^{(p)}=S_{z}^{(n)}=0$ and labeled $|\psi_{00}\rangle$, we consider two terms with a symmetric distribution of protons in the lower and higher energy orbitals, of $(2,0)$ and $(0,2)$. For the corresponding neutron states, the distribution of neutron holes is $(0,2)$ and $(2,0)$,
@@
Figure 9: Relative errors in energy $\epsilon_{E}$ for the same FH simulations with interaction $U$, middle hopping $t_{m}$ pairs and ”cuts” as shown in Fig. 6. Horizontal lines mark the errors $\epsilon_{E}$ determined by the corresponding Schmidt decompositions of the given layer (lower blue line). These errors can also be seen to decrease substantially as $t_{m}$ and entanglement diminish.
@@
In Appendix C we show other more technical improvements in the optimization of 60Ti, which go beyond splitting the circuits into halves and fourths and which allow to speed up the simulation.
Appendix B Further details on the convergence of EDEF in the nuclear shell model
@@
In the main text we showed the convergence of the infidelity in our examples as the most appropriate metric to test the method. The training, however, is based on minimizing the energy. For this reason, we also include here the relative errors in the energy, $\epsilon_{E}$, as the iterations of the optimization progress. Namely, in Fig. 9 we show the counterpart to Fig. 6, and in Fig. 10 the counterpart to Fig. 7. We can see that in both physical systems the behaviour of this error is qualitatively the same as that of infidelity, commented on the main text. In this case, relative energy errors given by the corresponding Schmidt decompositions, marked with horizontal lines, do not define a lower bound. In fact, for the 28Ne 1-cut EDEF simulation, the converged $\epsilon_{E}$ is slighlty lower than its corresponding Schmidt decomposition $\epsilon_{E}$.
@@
Figure 10: Relative errors in ground state energy for 28Ne (top panel) and 60Ti (bottom pane) as a function of the number of ADAPT-VQE iterations for 0-, 1-, and 2-cuts EDEF simulations.
@@
We also extend the amount of detail in the CNOT counts, $N_{\rm{CNOT}}$, by taking into account the effect of each subcircuit. Fig. 11 shows the number of CNOTs for regular ADAPT-VQE (solid black line on the top of the plot), together with $N_{\rm{CNOT}}$ for each circuit in the simulation with one layer of entanglement forging. $N_{\rm{CNOT}}$ for all 1-cut EDEF circuits are notably lower throughout all the evolution. Within single-cut EDEF lines we distinguish two-scales, $N_{\rm{CNOT}}\simeq 2\times 10^{3}$ for subcircuits simulating the first product state, with largest singular value, and $N_{\rm{CNOT}}\lesssim 10^{3}$ for the rest of subcircuit.
@@
Figure 11: CNOT count in each circuit for the simulation of 60Ti as function of ADAPT-VQE iterations for one cut.
In Fig. 12 we show similarly the CNOT counts for all subcircuit with 2-cut EDEF. The leading circuit is the neutron state with $S_{z}=0$.
Figure 12: CNOT count in each circuit for the simulation of 60Ti as function of ADAPT-VQE iterations for one cut.
@@
The simulations of 60Ti show similar results to 28Ne in terms of number of CNOTs. Specifically, 60Ti with one cut also shows a large difference between $N_{\rm{CNOT}}$ for $S_{z}=0$ and $S_{z}=1,2$ circuits. The former are of the order of $10^{3}$ for each proton and neutron circuits after 40 iterations, where ADAPT-VQE is not yet converged, see Fig. 13. This is similar to the neon isotopes, where ADAPT-VQE chooses in most of the iterations to apply an operator to one of the circuits corresponding to the first product state.
@@
Figure 13: CNOT count in each circuit for the simulation of 60Ti as function of ADAPT-VQE iterations for one cut.
@@
CNOT counts for 60Ti with two cuts are plotted in Fig. 14. The circuits with most CNOTs are much shallower than 1-cut ADAPT-VQE with the same number of iterations. In this case circuit with largest $N_{CNOT}$ is the one for neutrons with $S_{z}=0$ corresponding to the top subshells, which contains 660 CNOTs.
Figure 14: Number of CNOTs in each circuit for 60Ti as function of ADAPT-VQE iterations for two cuts.

### v2

@@ IV.3 Number of CNOT gates in EDEF NSM simulations
Figure 8 compares the number of CNOT gates, $N_{\rm CNOT}$, required in the NSM simulation of 28Ne and 60Ti in the case of regular ADAPT-VQE and EDEF with one or two layers. Since EDEF involves more than one circuit, we select the circuit with the maximum number of CNOT gates to plot Fig. 8. The top panel shows that, for 28Ne, the number of CNOTs in the $24$-qubit circuit (ADAPT-VQE, solid red line) increases drastically with the number of iterations. By the $100^{\rm th}$ iteration, more than $10^{4}$ CNOT gates are required. This number is however drastically reduced by one order of magnitude throughout the whole optimization when we employ EDEF with one layer, using $12$ qubits (1-cut EDEF, dashed blue line). Moreover, the $6$-qubit EDEF with two layers (2-cut EDEF, dotted-dashed yellow line) reduces the number of CNOTs by an additional order of magnitude. Consistently, the bottom panel of Fig. 8 indicates that for 60Ti the number of CNOTs is smaller in the 2-cut EDEF, which uses $10$-qubit circuits, than in the 1-cut EDEF, involving $20$ qubits. These results highlight that our EDEF approach, in addition to having better performance than the regular ADAPT-VQE as shown in Sec. IV.2, also requires substantially fewer quantum resources for these systems. As a consequence, the method is specially suited for near-future quantum device implementations.
@@
The reasons underlying the significant reduction of CNOT gates within the EDEF approach are twofold. First, each EDEF cut makes circuits considerably shallower because at each iteration an operator is added to a single circuit. This distributes the CNOT gates among all circuits, rather than accumulating them in a single one. Second, circuits with fewer qubits constrain the fermionic operators in the ADAPT-VQE pool. Because of the cuts, these operators are more local (eg connecting nearby qubits). This requires less CNOT gates in the corresponding Jordan-Wigner mapping, reducing the overall costs.
@@
Figure 8: Number of CNOT gates for the circuit with maximum number of them for the simulation of 28Ne (top) and 60Ti (bottom) as a function of the ADAPT-VQE iterations. The results compare the regular ADAPT-VQE (0 cuts, solid red line) with EDEF with one layer (1 cut, blue dashed line) and two layers (2 cuts, dashed-dotted yellow lines).
@@
Among all the EDEF subcircuits describing both 60Ti and 28Ne, the ones with most CNOT gates are those simulating the first, non-degenerate product state in the Schmidt decomposition, $\lambda_{0}|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle$ in Eq. (14). This is because the ADAPT-VQE algorithm finds, in most iterations, largest gradients for operators added to the first $\langle J_{z}\rangle=0$ proton-neutron product state. This is expected from the Schmidt decomposition, as the first singular value is much larger than the rest and should therefore contribute most to lowering the ground-state energy, at least in the initial iterations. Moreover, we also find that ADAPT-VQE chooses the same circuit for many contiguous iterations, showing plateaus that indicate a constant number of CNOTs in all but one circuit during various iterations. We provide more details on the specifics of the ADAPT-VQE optimization with EDEF in Appendix D, where we also break down the number of CNOT gates for each of the circuits that appear in the EDEF approach.
@@ V Summary and outlook
We present a novel procedure, entropy-driven entanglement forging (EDEF), that optimizes the solution of many-body problems in quantum computers using VQAs. Our approach is based on entanglement forging, which reduces the number of qubits and quantum gates needed to solve the problem. Crucially, in order to decide on the corresponding partitions and coefficients, we advocate for physical criteria associated to the entropy between partitions of the system. To this end, we explore systems that we expect to have low entropy partitions and employ the Schmidt decomposition as a tool to characterize them through their different quantum numbers. We build reference states that are linear combinations of smaller circuits with the quantum numbers associated to the Schmidt decomposition. Moreover, we exploit degeneracy in the lowest-lying Schmidt vectors to employ a very small number of subcircuits and coefficients, requiring fewer and more local resources.
@@
With this technique, we have successfully simulated ground states of FH lattices with various central hoppings and interactions, and of the isotopes 28Ne and 60Ti calculated with the NSM. In both cases, we compare optimizations using ADAPT-VQE with and without EDEF. For both types of systems, EDEF exploits the low entanglement between different partitions of the system. In the FH model, we can tune the entanglement between left and right parts of the lattice through a central hopping term with variable strength, displaying the power of the technique on different settings. In contrast, we exploit the recently observed small entanglement between protons and neutrons in the NSM to improve upon previous simulation attempts. This knowledge is combined with the singular-value degeneracy in the Schmidt decomposition of the ground states. Our EDEF simulations with one layer of EDEF – separating the system in two unique partitions – only need half as many qubits as the regular ADAPT-VQE, and they describe nuclei with an order of magnitude fewer CNOT gates. Nuclei simulated with two EDEF layers, with each bipartition additionally separated into two parts corresponding to higher- and lower-energy single-particle states, need a quarter of the regular ADAPT-VQE qubits and an additional order of magnitude fewer CNOTs.
@@
In terms of performance, in FH lattices, we find that EDEF converges better than regular ADAPT-VQE for weak central hopping. The approach is also more efficient in terms of iterations as long as the central hopping is close to the other hoppings, $t_{m}\approx t$. In the NSM, for 28Ne, one layer of EDEF using 8 circuits is more efficient than the regular ADAPT-VQE. Furthermore, EDEF allows us to simulate 60Ti, which is beyond our capabilities with the standard ADAPT-VQE. For this nucleus, the two layer EDEF reaches a lower infidelity than the 1-cut EDEF.
@@
In summary, EDEF is an approach ideally suited to accommodate the qubit and CNOT gate number limitations present in current intermediate-scale noisy devices. In particular, EDEF allows one to adjust algorithms to produce shallower circuits, mitigating the impact of errors. This reduction can be further improved by the adaptability of the trained parameters expected for a variational algorithm. Furthermore, by decreasing the number of qubits, EDEF allows one to use smaller devices with lower error rates. In principle, we have shown that these advantages can be exploited further by identifying low-entanglement partitions in successive EDEF layers. Beyond the first applications studied in this work, EDEF can be used to study other many-body systems for which low-entropy subsystems are identified. More generally, EDEF can potentially be used to optimize other types of VQAs for different applications, including for instance quantum machine learning settings.
@@ Acknowledgements.
This work is financially supported by MCIN/AEI/10.13039/501100011033 from the following grants: PID2020-118758GB-I00, PID2021-127890NB-I00, PID2023-147112NB-C22, RYC-2017-22781 and RYC2018-026072 through the “Ramón y Cajal” program funded by FSE “El FSE invierte en tu futuro”, CNS2022-135529 and CNS2022-135716 funded by the “European Union NextGenerationEU/PRTR”, and CEX2019-000918-M to the “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences; and by the Generalitat de Catalunya, grants 2021SGR01095 and 2021SGR00907.
@@ A.1 First layer of entanglement forging
For the nuclear shell model and for all isotopes considered, with one layer of entanglement forging we separate the proton and neutrons degrees of freedom as shown in Fig. 4 and thus reduce by half the number of qubits of the simulating circuits. As observed in Fig. 5, the nuclear state can be very accurately approximated using the following truncation in the Schmidt decomposition
@@
With one circuit, for example, we simulate the state $|p_{-2}\rangle$, which corresponds to a superposition of Slater determinants containing only proton orbitals and with expected third component of total angular momentum $\langle p_{-2}|J_{z}|p_{-2}\rangle=-2$. We need to ensure these states are orthogonal. All states that have different $\langle J_{z}\rangle$ in each partition are already orthogonal, and will continue to be considering the operators in our pool. This is however not the case for the two states with $\langle J_{z}\rangle=0$, $|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle$ and $|p_{0}\rangle\otimes|n_{0}\rangle$. For these, we start with different (orthogonal) Slater determinants, $|\psi_{0}\rangle$, $|\tilde{\psi}_{0}\rangle$, and then apply the same unitaries to both states to keep them orthogonal, that is, $\langle\tilde{\psi}_{0}|e^{-i\theta T}e^{i\theta T}|\psi_{0}\rangle=\langle\tilde{\psi}_{0}|\psi_{0}\rangle=0$.
At each iteration the gradients are computed for each of the terms in Eqs. (13) and (14) separately, except for $|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle$ and $|p_{0}\rangle\otimes|n_{0}\rangle$, as they both have the same parameters. In this case $|\psi_{c}\rangle$ in Eq. (10) represents $\lambda_{0}|\tilde{p}_{0}\rangle\otimes|\tilde{n}_{0}\rangle+\lambda_{1}|p_{0}\rangle\otimes|n_{0}\rangle$. Once all the gradients are computed, the largest one determines which operator from the pool is chosen and in which circuit is implemented. To compute the energy, we extract the statevector from each circuit, make the corresponding tensor products between proton and neutron states, add them up as in Eq. (14), and compute the expected value of the Hamiltonian with the obtained statevector. The energy computation in an actual quantum computer could be implemented as proposed in [29], with the exception of matrix elements of the Hamiltonian involving different product states.
@@ A.2 Second layer of entanglement forging
Figure 9: Relative errors in energy $\epsilon_{E}$ for the same FH simulations with interaction $U$, middle hopping $t_{m}$ pairs and ”cuts” as shown in Fig. 6. Horizontal lines mark the errors $\epsilon_{E}$ determined by the corresponding Schmidt decompositions of the given layer (lower blue line). These errors decrease substantially as $t_{m}$, and thus entanglement, diminish from right to left.
To add only few more parameters to the ansatz we decompose each of the previous statevectors with well defined spin and isospin into only two terms with different particle distributions. For the first term in Eq. (14), with $J_{z}^{(p)}=J_{z}^{(n)}=0$ and labeled $|\psi_{00}\rangle$, we consider two terms with a symmetric distribution of protons in the lower and higher energy orbitals, of $(2,0)$ and $(0,2)$. For the corresponding neutron states, the distribution of neutron holes is $(0,2)$ and $(2,0)$,
@@
In Appendix D we show other more technical improvements in the optimization of 60Ti, which go beyond splitting the circuits into halves and fourths and which allow to speed up the simulation.
Appendix B Further details on the convergence of EDEF
@@
In the main text, we show the convergence of the infidelity in two physical systems. This is an appropriate metric to test the method since the infidelity ties directly to the Schmidt decomposition. The training of the circuit, however, is based on an energy minimisation. For this reason, we also include and discuss here the relative errors in the energy, $\epsilon_{E}$, as the iterations of the optimization progress. Namely, in Fig. 9 we show the counterpart to Fig. 6 for the FH in terms of energy. We also show in Fig. 10 the counterpart to Fig. 7 for the NSM. The notation of the figures follows that of the main text.
The two figures indicate that, in both systems, the behaviour of the relative energy error is qualitatively similar to that of the infidelity, already described in the main text. For the FH model, in particular, Fig. 9 is also quantitatively similar to Fig. 6. We find that EDEF outperforms standard ADAPT-VQE with the system settings of panel (d), with mild entanglement between the partitions. In all other cases, the EDEF minimisation saturates an energy bound calculated indirectly from the Schmidt decomposition within about $15$ iterations. This bound is not however a strict lower bound for the EDEF simulations. Indeed, panels (c) and (e) show examples where the bound is mildly improved by EDEF.
@@
Figure 10: Relative errors in ground state energy for 28Ne (top) and 60Ti (bottom) as a function of the number of ADAPT-VQE iterations for 0-, 1-, and 2-cuts EDEF simulations.
For the NSM, the relative energy analysis in Fig. 10 also shows qualitative agreement to the infidelity results presented in Fig. 7 for 28Ne and 60Ti. Here, however, the quantitative discrepancies are more visible. First, the ADAPT-VQE energies for 28Ne decrease steadily with iterations, unlike the corresponding infidelities of Fig. 7, which increase at times. The rate of energy decrease for the full ADAPT-VQE simulation, however, is relatively slow: we require $100$ iterations to achieve a relative energy error of $\epsilon_{E}\approx 6\times 10^{-3}$. In contrast, the 1-cut EDEF reaches a relative energy error which is about an order of magnitude smaller with only $80$ iterations (and far fewer resources). We stress again that relative energy errors given by the corresponding Schmidt decompositions, marked with horizontal lines, do not define a lower bound. In fact, for the 28Ne 1-cut EDEF simulation, the converged $\epsilon_{E}$ is slightly lower than its corresponding Schmidt decomposition $\epsilon_{E}$. In contrast, the 2-cut EDEF simulation reaches an error of only $\epsilon_{E}\approx 10^{-2}$, but does so within less than $20$ iterations. This error is somewhat larger than the bound suggested by the Schmidt decomposition, a result that is commensurate with the infidelity analysis of Fig. 7. We take this as an indication of the limitation of the 2-cut approach, that cannot capture the residual entanglement of the second cut.
The results for 60Ti presented in the bottom panel of Fig. 10 show a similar picture, although at a different level of quality (note the difference in $y-$axis between the two panels). As discussed in the main text, memory limitations preclude us from simulating the statevector for this isotope with full ADAPT-VQE. The 1- and 2-cut EDEF simulations start relatively far away from the final result (with $\epsilon_{E}\approx 1$). They subsequently decrease in energy steadily, at a rate that is comparable for the two simulations. This is at odds with the infidelity results of Fig. 7, which indicates that the 2-cut infidelity decreases much faster than the 1-cut simulation. We note however that the minimisation process is such that the infidelity and the energy lie relatively far from the Schmidt bounds.
Figure 11: CNOT count in each circuit for the simulation of 28Ne as function of iteration number. Full ADAPT-VQE results (black solid line) are compared to the $N_{\text{CNOT}}$ for each subcircuit of 1-cut EDEF simulation. We distinguish circuits with different third component $J_{z}$, see Eq. (19).
@@
We also extend the amount of detail in the CNOT counts, $N_{\rm{CNOT}}$, by taking into account the effect of each subcircuit for 1-cut (Fig. 11) and 2-cut (Fig. 12) EDEF. We perform this analysis for the NSM only in the interest of brevity. Figure 11 shows the number of CNOTs for regular ADAPT-VQE (solid black line), together with $N_{\rm{CNOT}}$ for each of the circuits in the simulation of 28Ne with one layer of entanglement forging. First of all, we stress the fact that $N_{\rm{CNOT}}$ for all 1-cut EDEF circuits are notably lower throughout all the evolution. We find substantially different $N_{\rm{CNOT}}$ values depending on the state represented by the different circuits. The subcircuits associated to the first product states, $|p_{0}\rangle$ and $|n_{0}\rangle$, and the equivalent circuits for $|\tilde{p}_{0}\rangle$ and $|\tilde{n}_{0}\rangle$, that start from a different state but are otherwise identical, have a much larger $N_{\rm{CNOT}}$ count, as shown in the red and puple solid lines in Fig. 11. In contrast, the states $|p_{1}\rangle$, $|p_{2}\rangle$, $|n_{1}\rangle$ and $|n_{2}\rangle$ are represented by much shallower circuits. In particular, these circuits are never selected by the VQA before the tenth iteration. The evolution of the CNOT count for these states is much more staggered. Each time an operator in the circuit is selected, the $N_{\rm{CNOT}}$ count raises substantially with a single, relatively large step. The different staggering for each subcircuit suggests that each circuit is only picked up in the minimisation process at different stages. Clearly, in the initial stages of the minimisation the dominant product states with $\langle J_{z}\rangle=0$ are picked up, whereas states with lower Schmidt coefficients are only selected later in the minimisation process. In terms of total counts, within single-cut EDEF, we distinguish two scales: $N_{\rm{CNOT}}\simeq 2\times 10^{3}$ for subcircuits simulating the first product state, with largest singular value, and $N_{\rm{CNOT}}\lesssim 10^{3}$ for the rest of subcircuit.
@@
Figure 12: CNOT count in each circuit for the simulation of 28Ne as function of iteration number. We show $N_{\text{CNOT}}$ for each subcircuit of a 2-cut EDEF simulation.
We provide a similar analysis for the CNOT counts for all subcircuits with 2-cut EDEF for 28Ne in Fig. 12. Here, we focus on the eight subcircuits which contain CNOT gates with different Schmidt coefficients, belonging to proton and neutron partitions at the top or bottom of the single-particle spectrum. The number of CNOT gates in these 2-cut simulations is much smaller than the corresponding values of 1-cut simulations. The maximum number is well below $400$ across all the iterations, and corresponds to the dominant state in the neutron partition with single-particle energies at the top of the spectrum (eg $1s_{1/2}$ and $0d_{3/2}$). This leading circuit has $\langle J_{z}\rangle=0$ and is shown in a purple dotted line. The corresponding proton subcircuit (dash-dotted orange line) is the second most CNOT intensive circuit, reaching $N_{\rm CNOT}\approx 200$ at the end of the simulation. Neutron and proton bottom partitions (with $0d_{5/2}$ configurations) end up the minimization with about $N_{\rm CNOT}\approx 100-150$. The remaining circuits correspond to smaller Schmidt coefficients and have less than $100$ CNOT gates by the time the simulation finishes. Overall, this bodes well with the idea that the subcircuits with largest Schmidt coefficients are picked up first, and more often, in the minimisation process.
@@
The simulations of 60Ti show similar results to 28Ne in terms of number of CNOTs for each subcircuit. We show in Fig. 13 the evolution of different subcircuit $N_{\rm CNOT}$ counts for the 1-cut simulation of 60Ti. 60Ti with one cut also shows a large difference between $N_{\rm{CNOT}}$ for $\langle J_{z}\rangle=0$ and $\langle J_{z}\rangle=1,2$ circuits. This is similar to the neon isotopes, where ADAPT-VQE chooses in most of the iterations to apply an operator to one of the circuits corresponding to the first product state. In fact, an operator is not applied to the second product states until the $20^{\text{th}}$ iteration. By the point that the simulation is converged, at iteration $50$, the proton and neutron circuits for the first product state have about $10^{3}$ CNOT gates each. For this neutron-rich isotope, however, the circuits associated to neutrons with $\langle J_{z}\rangle=1$ and $2$ have a similar number of gates, in spite of the relatively smaller Schmidt coefficient. The remaining proton subcircuits require about $500$ gates each.
@@
Figure 13: CNOT count in each circuit for the simulation of 60Ti as function of iteration number. We show $N_{\text{CNOT}}$ for each subcircuit of a 1-cut EDEF simulation.
Figure 14: CNOT count in each circuit for the simulation of 60Ti as function of iteration number. We show $N_{\text{CNOT}}$ for each subcircuit of a 2-cut EDEF simulation.
@@
CNOT counts for 60Ti with two cuts are plotted in Fig. 14. The circuits with most CNOTs are much shallower than 1-cut ADAPT-VQE with the same number of iterations. In this case, the circuit with the largest $N_{CNOT}$ is the one for neutrons with $\langle J_{z}\rangle=0$ corresponding to the top subshells, which contains $660$ CNOTs. This is followed by the corresponding $\langle J_{z}\rangle=0$ proton circuit for the bottom partition, with over $400$ gates. Whereas for 28Ne, the $\langle J_{z}\rangle=0$ proton state for the top single-particle partition was the second subcircuit in terms of CNOTs, for 60Ti the same state ranks sixth in terms of CNOT counts. This is indicative of the individual complex pathways in the minimization process for different isotopes.
Appendix C Results for a different isotope